# Simultaneous spatial-parametric collocation approximation for parametric PDEs with log-normal random inputs

### v4

@@ 2.3 Multi-level least squares sampling algorithms
$\begin{split}\sup_{v\in B_{X}^{\alpha}}\|v-\bar{{\mathcal{S}}}^{X}_{n}v\|_{L_{2}(U,X;\mu)}\ \leq\ \begin{cases}C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ \alpha<1/q_{2},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha=1/q_{2},\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>1/q_{2}.\end{cases}\end{split}$
@@
Notice that if $\|{\boldsymbol{\sigma}}^{-1}\|_{\ell_{q}({\mathbb{F}})}<\infty$ for some $0<q\leq 2$, then for every $v\in B_{X,{\boldsymbol{\sigma}}}({\mathbb{U}}^{\infty})$, the series (2.37) converges absolutely and unconditionally in $L_{2}({\mathbb{U}}^{\infty},X;\gamma)$ to $v$ (see [22, Lemma 3.1] for the case $\mathbb{U}$$^{\infty}$$\mathbb{R}$$^{\infty}$, the case ${\mathbb{U}}^{\infty}={\mathbb{I}}^{\infty}$ can be proven by the same arguments). Hence, we can reorder the countable set $\mathbb{F}$} as ${\mathbb{F}}=({\boldsymbol{s}}_{j})_{j\in{\mathbb{N}}}$ so that the sequence ${\boldsymbol{\sigma}}=(\sigma_{{\boldsymbol{s}}_{j}})_{j\in{\mathbb{N}}}$ is non-decreasing. Put $U:={\mathbb{U}}^{\infty}$, $\sigma_{j}:=\sigma_{{\boldsymbol{s}}_{j}}$, $\varphi_{j}:=\phi_{{\boldsymbol{s}}_{j}}$ and $v_{j}:=v_{{\boldsymbol{s}}_{j}}$. Then $B_{X,{\boldsymbol{\sigma}}}({\mathbb{U}}^{\infty})$ can be seen as the set $B_{X,{\boldsymbol{\sigma}}}$ of all functions $v\in L_{2}(U,X;\gamma))$ represented by the series
@@
Theorem 2.2 for the space $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$ is read as
@@
$\begin{split}\sup_{v\in B_{X}^{\alpha}}\|v-\bar{{\mathcal{S}}}^{X}_{n}v\|_{L_{2}({\mathbb{R}}^{\infty},X;\gamma)}\ \leq\ \begin{cases}C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ \alpha<1/q_{2},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha=1/q_{2},\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>1/q_{2}.\end{cases}\end{split}$
@@ 3 Applications to holomorphic functions
where q:=p/(1p)q:=p/(1-p), with $|{\boldsymbol{s}}^{\prime}|_{\infty}:=\sup_{j\in{\mathbb{N}}}s^{\prime}_{j}$ the set ${\boldsymbol{\sigma}}:=(\sigma_{\boldsymbol{s}})_{{\boldsymbol{s}}\in{\mathbb{F}}}$ is defined by
@@
$M=\delta C_{{\boldsymbol{b}},\xi,\eta}$ with some positive constant $C_{{\boldsymbol{b}},\xi,\eta}$, and $N=K_{{\boldsymbol{b}},\xi,\eta}$.
@@
$\alpha>0$, $\eta\in{\mathbb{N}}$, $0<p_{1}\leq p_{2}<1$, $0<p_{1}<2/3$, ${\boldsymbol{b}}_{i}\in\ell_{p_{1}}({\mathbb{N}})$, $i=1,2$, are given sequences, $\xi>0$, $\delta>0$. The sequences ${\boldsymbol{\rho}}_{i}=(\rho_{i,j})_{j\in{\mathbb{N}}}$, $i=1,2$, are defined by
$\rho_{i,j}:=b_{i,j}^{p_{i}-1}\frac{\xi}{4\sqrt{\eta!}\left\|{{\boldsymbol{b}}_{i}}\right\|_{\ell_{p_{j}}({\mathbb{N}})}},\ j\in{\mathbb{N}},$ (3.2)
${\mathcal{V}}=(V_{k})_{k\in{\mathbb{N}}_{0}}$ is a given sequence of subspaces $V_{k}\subset L_{2}({\mathbb{R}}^{\infty},X;\gamma)$ of dimension $2^{k}$, and ${\mathcal{P}}=(P_{k})_{k\in{\mathbb{N}}_{0}}$ a given sequence of linear uniformly bounded projectors from $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$ onto $V_{k}$ such that
@@
$\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}(\sigma_{1,{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X})^{2}\right)^{1/2}\ \leq M,\ \ \text{with}\ \ \left\|{{\boldsymbol{\sigma}}_{1}^{-1}}\right\|_{\ell_{q}({\mathbb{F}})}\leq N<\infty,$
@@
$\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}(\sigma_{2,{\boldsymbol{s}}}\|(v-P_{k}v)_{\boldsymbol{s}}\|_{X})^{2}\right)^{1/2}\ \leq M2^{-\alpha k}\ <\infty,\ \ \text{with}\ \ \left\|{{\boldsymbol{\sigma}}_{2}^{-1}}\right\|_{\ell_{q}({\mathbb{F}})}\leq N<\infty,$
@@
$\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}(\sigma_{1,{\boldsymbol{s}}}\|{\color[rgb]{0,0,0}{(P_{m}v)_{\boldsymbol{s}}}}\|_{X})^{2}\right)^{1/2}\ \leq\ K_{{\boldsymbol{b}},\xi,\eta}M,\ \ \left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}(\sigma_{2,{\boldsymbol{s}}}\|{\color[rgb]{0,0,0}{(P_{m}v-P_{k}(P_{m}v))_{\boldsymbol{s}}}}\|_{X})^{2}\right)^{1/2}\ \leq\ K_{{\boldsymbol{b}},\xi,\eta}M2^{-\alpha k}.$
@@
Let $v$ satisfy Assumption 3.1. Let $q_{i}:=p_{i}/(1-p_{i})$ for $i=1,2$, and $\varepsilon$ be any positive number. Let the number $\beta$ be defined as in (2.25). Let $(\xi_{n})_{n\in{\mathbb{N}}}$ be a sequence of increasing positive numbers $\xi_{n}$ satisfying the condition (2.32). Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\gamma)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{n}:={\mathcal{S}}_{\lceil n/\log n\rceil}^{X}$ defined as in (2.29)–(2.31),
@@
$\begin{split}\|v-\bar{{\mathcal{S}}}^{X}_{n}v\|_{L_{2}({\mathbb{R}}^{\infty},X;\gamma)}\ \leq\ \begin{cases}C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ \alpha<1/q_{2},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha=1/q_{2},\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>1/q_{2}.\end{cases}\end{split}$
@@ 4 Multi-level sparse-grid interpolation algorithms
In this section, we construct fully discrete multi-level sparse-grid sampling algorithms for the $({\boldsymbol{b}},\xi,\delta,X)$-holomorphic functions satisfying Assumption 3.1 except the item (iv) , based on GPC Lagrange-Hermite interpolation, and prove convergence rates of the approximation by them. It terns out that in the case $\alpha\leq 1/q_{2}-1/2$, these sampling algorithms give a convergence rate better than the extended least squares sampling algorithm in Theorem 3.2. The construction and techniques used in this section are a modification of those in [23] .
@@ 5 Applications to parametric elliptic PDEs
In this section, we prove that the parametric solution $u$ to the parametric elliptic PDEs (1.1) on a bounded polygonal domain with log-normal inputs (1.3) is a $({\boldsymbol{b}}_{j},\xi,\delta,V)$-holomorphic function on $\mathbb{R}$$^{\infty}$ satisfying Assumption 3.1 with ${\boldsymbol{b}}_{j}$ as in (5.4), $j=1,2$. This allows to establish convergence rates of a fully discrete multi-level collocation algorithm $S_{n}$. The spatial components in $S_{n}$ are based on finite element Lagrange interpolations associated with triangulations of the spatial domain $D$. While the parametric components are based on Hermite-Lagrange GPC interpolation as in Theorem 5.3 in the case of small spatial regularity $\alpha\leq 1/p_{2}-3/2$, and on extended least squares sampling algorithms as in Theorem 3.2 in the cases of higher spatial regularity $\alpha>1/p_{2}-3/2$.
@@
$b_{1,j}:=\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{L_{\infty}(D)}}},\quad b_{2,j}:=\max\big\{\left\|{\psi_{j}}\right\|_{W^{1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{{\mathcal{W}}_{\infty}^{r-1}(D)}\big\}$ (5.4)
and ${\boldsymbol{b}}_{1}:=(b_{1,j})_{j\in{\mathbb{N}}}$, ${\boldsymbol{b}}_{2}:=(b_{2,j})_{j\in{\mathbb{N}}}$.
@@
Let $D\subset{\mathbb{R}}^{2}$ be a bounded polygonal domain and $r\in{\mathbb{N}}$, $r\geq 2$. Let $\alpha=\frac{r-1}{2}$. Let $\psi_{k}\in W^{1}_{\infty}(D)\cap{\mathcal{W}}_{\infty}^{r-1}(D)$, $k\in{\mathbb{N}}$, and $f\in{\mathcal{K}}_{\varkappa-1}^{r-2}(D)$. Let $0<p_{1},p_{2}<1$. For $i=1,2$, let the sequences ${\boldsymbol{b}}_{i}:=(b_{i,j})_{j\in{\mathbb{N}}}$ be defined as in (5.4) with ${\boldsymbol{b}}_{i}\in{\color[rgb]{0,0,0}{\ell_{p_{i}}}}({\mathbb{N}})$. Then there exist $\xi>0$ and $\delta>0$ such that for the parametric solution to (5.1) with inputs (1.3)
@@
Step 2. Concerning the claim (2), by assumption, $b_{1,j}=\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{L_{\infty}(D)}}}$ satisfies ${\boldsymbol{b}}_{1}=(b_{1,j})_{j\in{\mathbb{N}}}\in{\color[rgb]{0,0,0}{\ell_{p_{1}}({\mathbb{N}})\subseteq\ell_{1}({\mathbb{N}})}}$, which corresponds to the assumption (iv) of Theorem [25, Theorem 4.11]. It remains to verify the assumptions (i), (ii) and (iii) of [25, Theorem 4.11] for $P_{k}u$, $k\in{\mathbb{N}}_{0}$.
@@
$b_{2,j}:=\max\big\{\left\|{\psi_{j}}\right\|_{{\mathcal{W}}^{s-1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{W^{1}_{\infty}(D)}}}\big\},$
@@
${\boldsymbol{b}}_{2}=(b_{2,j})_{j\in{\mathbb{N}}}\in{\color[rgb]{0,0,0}{\ell_{p_{2}}({\mathbb{N}})\hookrightarrow\ell_{1}({\mathbb{N}})}},$
@@
$D\subset{\mathbb{R}}^{2}$ is a bounded polygonal domain and $r\in{\mathbb{N}}$, $r\geq 2$; $f\in{\mathcal{K}}_{\varkappa-1}^{r-2}(D)$ and $\psi_{k}\in W^{1}_{\infty}(D)\cap{\mathcal{W}}_{\infty}^{r-1}(D)$, $k\in{\mathbb{N}}$. For $i=1,2$, the sequences ${\boldsymbol{b}}_{i}:=(b_{i,j})_{j\in{\mathbb{N}}}$ defined as in (5.4) satisfy the condition ${\boldsymbol{b}}_{i}\in{\color[rgb]{0,0,0}{\ell_{p_{i}}}}({\mathbb{N}})$ with $0<p_{1}\leq p_{2}<1$ and $p_{1}<2/3$.
@@
$\alpha:=\frac{r-1}{2},\quad\beta:=\left(\frac{1}{p_{1}}-1\right)\frac{\alpha}{\alpha+\delta},\quad\delta:=\frac{1}{p_{1}}-\frac{1}{p_{2}}.$ (5.9)
@@
$\begin{split}\big\|u-\bar{{\mathcal{S}}}^{V}_{n}u\big\|_{L_{2}({\mathbb{R}}^{\infty},V;\gamma)}\ \leq\ \begin{cases}C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ \alpha<1/q_{2},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha=1/q_{2},\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>1/q_{2},\end{cases}\end{split}$
where the constants $C$ and $C_{\varepsilon}$ is independent of $n$.
@@
$\alpha:=\frac{r-1}{2},\quad\beta:=\left(\frac{1}{p_{1}}-\frac{3}{2}\right)\frac{\alpha}{\alpha+\delta},\quad\delta:=\frac{1}{p_{1}}-\frac{1}{p_{2}}.$ (5.11)
@@
$\big|\operatorname{SamplePts}{\mathcal{I}}_{\xi_{n}}\big|\ \leq\ n,$
@@
$\begin{split}\big\|u-{\mathcal{I}}_{\xi_{n}}u\big\|_{L_{2}({\mathbb{R}}^{\infty},V;\gamma)}\ \leq\ C\begin{cases}n^{-\alpha}\quad&{\rm if}\ \alpha\leq 1/p_{2}-3/2,\\ n^{-\beta}\quad&{\rm if}\ \alpha>1/p_{2}-3/2,\end{cases}\end{split}$

### v5

@@ 2.3 Multi-level least squares sampling algorithms
$\begin{split}\sup_{v\in B_{X}^{\alpha}}\|v-\bar{{\mathcal{S}}}^{X}_{n}v\|_{L_{2}(U,X;\mu)}\ \leq\ \begin{cases}C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ \alpha<1/q_{2},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha=1/q_{2},\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>1/q_{2},\end{cases}\end{split}$
where the positive constant $C$ depends only on $\alpha,q_{1},q_{2}$ and $K$; the positive constants $C_{\varepsilon}$ depends only on $\alpha,q_{1},q_{2},K,$ and, explicitly, on $\varepsilon$.
Remark 2.3.
We remark on how the constants in Theorems 2.1 and 2.2 depend on the relevant parameters. The dependence of the constants $C_{\tau}$ in the cases $\alpha<1/q_{2}$ and $\alpha=1/q_{2}$ in Theorem 2.1 are given explicitly in (2.37) and (2.38), respectively. In contrast, the constant $C$ in the case $\alpha>1/q_{2}$ depends on $\alpha,q_{1},q_{2}$ and $K$ only since $\tau:=\alpha/(\alpha-\beta)$. In Theorem 2.2, for the cases $\alpha<1/q_{2}$ and $\alpha=1/q_{2}$, the positive constants $C_{\varepsilon}$ depend only on $\alpha,q_{1},q_{2},K,$ and, explicitly, on $\varepsilon$.
Similar remarks regarding the constants in Theorems 6.2 and 6.3 apply.
@@
Notice that if $\|{\boldsymbol{\sigma}}^{-1}\|_{\ell_{q}({\mathbb{F}})}<\infty$ for some $0<q\leq 2$, then for every $v\in B_{X,{\boldsymbol{\sigma}}}$, the series (2.39) converges absolutely and unconditionally in $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$ to $v$ (see [22, Lemma 3.1]). Hence, we can reorder the countable set $\mathbb{F}$} as ${\mathbb{F}}=({\boldsymbol{s}}_{j})_{j\in{\mathbb{N}}}$ so that the sequence ${\boldsymbol{\sigma}}=(\sigma_{{\boldsymbol{s}}_{j}})_{j\in{\mathbb{N}}}$ is non-decreasing. Put $\sigma_{j}:=\sigma_{{\boldsymbol{s}}_{j}}$, $\varphi_{j}:=\phi_{{\boldsymbol{s}}_{j}}$ and $v_{j}:=v_{{\boldsymbol{s}}_{j}}$. Then $B_{X,{\boldsymbol{\sigma}}}$ can be seen as the set of all functions $v\in L_{2}(U,X;\gamma))$ for $U:={\mathbb{R}}^{\infty}$, represented by the series
@@
This means that the pointwise evaluations $f({\boldsymbol{y}})$ are well-defined for every $\boldsymbol{y}$} belonging the full measure subset $U_{0}$ of $U$. Let $\left(\psi_{k}\right)_{k\in{\mathbb{N}}}$ be an orthonormal basis of $X$. If $v\in H_{X,{\boldsymbol{\sigma}}}$, then $\langle v,\psi_{k}\rangle_{X}\in H_{{\mathbb{C}},{\boldsymbol{\sigma}}}$ for every $k$. Consequently, the pointwise evaluations $\langle v({\boldsymbol{y}}),\psi_{k}\rangle_{X}$ are well-defined for every ${\boldsymbol{y}}\in U_{0}$. This implies that the pointwise evaluations $v({\boldsymbol{y}})$ are also well-defined for every ${\boldsymbol{y}}\in U_{0}$.
@@
Let ${\mathbb{R}}^{\infty}_{0}\subset{\mathbb{R}}^{\infty}$ be a set of full measure such that if $v\in H_{X}^{\alpha}$, the pointwise evaluations $v({\boldsymbol{y}})$ are also well-defined for every ${\boldsymbol{y}}\in{\mathbb{R}}^{\infty}_{0}$. Theorem 2.2 for the space $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$ is read as
@@
$\sup_{v\in B_{X}^{\alpha}}\|v-\bar{{\mathcal{S}}}^{X}_{n}v\|_{L_{2}({\mathbb{R}}^{\infty},X;\gamma)}\ \leq\ \begin{cases}C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ \alpha<1/q_{2},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha=1/q_{2},\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>1/q_{2},\end{cases}$ (2.40)
where the positive constant $C$ depends only on $\alpha,q_{1},q_{2}$ and $K$; the positive constants $C_{\varepsilon}$ depend only on $\alpha,q_{1},q_{2},K,$ and, explicitly, on $\varepsilon$.
@@ 3 Applications to holomorphic functions
where q:=2p/(2p)q:=2p/(2-p), the set ${\boldsymbol{\sigma}}:=(\sigma_{\boldsymbol{s}})_{{\boldsymbol{s}}\in{\mathbb{F}}}$ is defined by
@@
with $|{\boldsymbol{s}}^{\prime}|_{\infty}:=\sup_{j\in{\mathbb{N}}}s^{\prime}_{j}$, and $M=\delta C_{{\boldsymbol{b}},\xi,\eta}$ with some positive constant $C_{{\boldsymbol{b}},\xi,\eta}$, and $N=K_{{\boldsymbol{b}},\xi,\eta}$.
@@
$\alpha>0$, $\eta\in{\mathbb{N}}$, $0<p_{1}\leq p_{2}<1$, ${\boldsymbol{b}}_{i}\in\ell_{p_{1}}({\mathbb{N}})$, $i=1,2$, are given sequences, $\xi>0$, $\delta>0$. The sequences ${\boldsymbol{\rho}}_{i}=(\rho_{i;j})_{j\in{\mathbb{N}}}$, $i=1,2$, are defined by
$\rho_{i;j}:=b_{i;j}^{p_{i}-1}\frac{\xi}{4\sqrt{\eta!}\left\|{{\boldsymbol{b}}_{i}}\right\|_{\ell_{p_{j}}({\mathbb{N}})}},\ j\in{\mathbb{N}},$ (3.2)
$(V_{k})_{k\in{\mathbb{N}}_{0}}$ is a given sequence of subspaces $V_{k}\subset L_{2}({\mathbb{R}}^{\infty},X;\gamma)$ of dimension $2^{k}$, and ${\mathcal{P}}=(P_{k})_{k\in{\mathbb{N}}_{0}}$ a given sequence of linear uniformly bounded projectors from $L_{2}({\mathbb{R}}^{\infty},X;\gamma)$ onto $V_{k}$ such that
@@
$\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}(\sigma_{1;{\boldsymbol{s}}}\|v_{\boldsymbol{s}}\|_{X})^{2}\right)^{1/2}\ \leq M,\ \ \text{with}\ \ \left\|{{\boldsymbol{\sigma}}_{1}^{-1}}\right\|_{\ell_{q}({\mathbb{F}})}\leq N<\infty,$
@@
$\left(\sum_{{\boldsymbol{s}}\in{\mathbb{F}}}(\sigma_{2;{\boldsymbol{s}}}\|(v-P_{k}v)_{\boldsymbol{s}}\|_{X})^{2}\right)^{1/2}\ \leq M2^{-\alpha k}\ <\infty,\ \ \text{with}\ \ \left\|{{\boldsymbol{\sigma}}_{2}^{-1}}\right\|_{\ell_{q}({\mathbb{F}})}\leq N<\infty,$
@@
[…]
@@
Let $v$ satisfy Assumption 3.1. Let $\varepsilon$ be any fixed positive number. Let the number $\beta$ be defined as in (2.25). Let $(\xi_{n})_{n\in{\mathbb{N}}}$ be a sequence of increasing positive numbers $\xi_{n}$ satisfying the condition (2.32). Then for all $n\geq 2$ there exist ${\boldsymbol{y}}_{k,1},...,{\boldsymbol{y}}_{k,n_{k}}\in{\color[rgb]{0,0,0}{{\mathbb{R}}^{\infty}_{0}}}$ and $h_{k,1},...,h_{k,n_{k}}\in L_{2}({\mathbb{R}}^{\infty},{\mathbb{C}};\gamma)$, $k=0,...,k_{\lceil n/\log n\rceil}$, such that for the operator $\bar{{\mathcal{S}}}^{X}_{n}:={\mathcal{S}}_{\lceil n/\log n\rceil}^{X}$ defined as in (2.29)–(2.31),
@@
$\begin{split}\|v-\bar{{\mathcal{S}}}^{X}_{n}v\|_{L_{2}({\mathbb{R}}^{\infty},X;\gamma)}\ \leq\ \begin{cases}C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ \alpha<{\color[rgb]{0,0,0}{1/p_{2}-1/2}},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha={\color[rgb]{0,0,0}{1/p_{2}-1/2}},\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>{\color[rgb]{0,0,0}{1/p_{2}-1/2}},\end{cases}\end{split}$
where the positive constant $C$ depends only on $\alpha,q_{1},q_{2}$ and $K$; the positive constants $C_{\varepsilon}$ depend only on $\alpha,q_{1},q_{2},K,$ and, explicitly, on $\varepsilon$.
@@ 4 Multi-level sparse-grid interpolation algorithms
In this section, we construct fully discrete multi-level sparse-grid sampling algorithms for the $({\boldsymbol{b}},\xi,\delta,X)$-holomorphic functions satisfying Assumption 3.1 except the item (iv) , based on GPC Lagrange-Hermite interpolation, and prove convergence rates of the approximation by them. It turns out that in the case $\alpha\leq 1/q_{2}-1/2$, these sampling algorithms give a convergence rate better than the extended least squares sampling algorithm in Theorem 3.2. The construction and techniques used in this section are a modification of those in [23] .
@@ 5 Applications to parametric elliptic PDEs
In this section, we prove that the parametric solution $u$ to the parametric elliptic PDEs (1.1) on a bounded polygonal domain with log-normal inputs (1.3) is a $({\boldsymbol{b}}_{j},\xi,\delta,V)$-holomorphic function on $\mathbb{R}$$^{\infty}$ satisfying Assumption 3.1 with ${\boldsymbol{b}}_{j}$ as in (5.4), $j=1,2$. This allows to establish convergence rates of a fully discrete multi-level collocation algorithm $S_{n}$. The spatial components in $S_{n}$ are based on finite element Lagrange interpolations associated with triangulations of the spatial domain $D$. While the parametric components are based on Hermite-Lagrange GPC interpolation as in Theorem 5.3 in the case of small spatial regularity $\alpha\leq 1/p_{2}-1$ , and on extended least squares sampling algorithms as in Theorem 3.2 in the cases of higher spatial regularity $\alpha>1/p_{2}-1$ .
@@
$b_{1;j}:=\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{L_{\infty}(D)}}},\quad b_{2;j}:=\max\big\{\left\|{\psi_{j}}\right\|_{W^{1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{{\mathcal{W}}_{\infty}^{r-1}(D)}\big\}$ (5.4)
and ${\boldsymbol{b}}_{1}:=(b_{1;j})_{j\in{\mathbb{N}}}$, ${\boldsymbol{b}}_{2}:=(b_{2;j})_{j\in{\mathbb{N}}}$.
@@
Let $D\subset{\mathbb{R}}^{2}$ be a bounded polygonal domain and $r\in{\mathbb{N}}$, $r\geq 2$. Let $\alpha=\frac{r-1}{2}$. Let $\psi_{k}\in W^{1}_{\infty}(D)\cap{\mathcal{W}}_{\infty}^{r-1}(D)$, $k\in{\mathbb{N}}$, and $f\in{\mathcal{K}}_{\varkappa-1}^{r-2}(D)$. Let $0<p_{1},p_{2}<1$. For $i=1,2$, let the sequences ${\boldsymbol{b}}_{i}:=(b_{i;j})_{j\in{\mathbb{N}}}$ be defined as in (5.4) with ${\boldsymbol{b}}_{i}\in{\color[rgb]{0,0,0}{\ell_{p_{i}}}}({\mathbb{N}})$. Then there exist $\xi>0$ and $\delta>0$ such that for the parametric solution to (5.1) with inputs (1.3)
@@
Step 2. Concerning the claim (2), by assumption, $b_{1;j}=\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{L_{\infty}(D)}}}$ satisfies ${\boldsymbol{b}}_{1}=(b_{1;j})_{j\in{\mathbb{N}}}\in{\color[rgb]{0,0,0}{\ell_{p_{1}}({\mathbb{N}})\subseteq\ell_{1}({\mathbb{N}})}}$, which corresponds to the assumption (iv) of Theorem [25, Theorem 4.11]. It remains to verify the assumptions (i), (ii) and (iii) of [25, Theorem 4.11] for $P_{k}u$, $k\in{\mathbb{N}}_{0}$.
@@
$b_{2;j}:=\max\big\{\left\|{\psi_{j}}\right\|_{{\mathcal{W}}^{s-1}_{\infty}(D)},\left\|{\psi_{j}}\right\|_{{\color[rgb]{0,0,0}{W^{1}_{\infty}(D)}}}\big\},$
@@
${\boldsymbol{b}}_{2}=(b_{2;j})_{j\in{\mathbb{N}}}\in{\color[rgb]{0,0,0}{\ell_{p_{2}}({\mathbb{N}})\hookrightarrow\ell_{1}({\mathbb{N}})}},$
@@
$D\subset{\mathbb{R}}^{2}$ is a bounded polygonal domain and $r\in{\mathbb{N}}$, $r\geq 2$; $f\in{\mathcal{K}}_{\varkappa-1}^{r-2}(D)$ and $\psi_{k}\in W^{1}_{\infty}(D)\cap{\mathcal{W}}_{\infty}^{r-1}(D)$, $k\in{\mathbb{N}}$. For $i=1,2$, the sequences ${\boldsymbol{b}}_{i}:=(b_{i,j})_{j\in{\mathbb{N}}}$ defined as in (5.4) satisfy the condition ${\boldsymbol{b}}_{i}\in{\color[rgb]{0,0,0}{\ell_{p_{i}}}}({\mathbb{N}})$ with $0<p_{1}\leq p_{2}<1$.
@@
$\alpha:=\frac{r-1}{2},\quad\beta:=\left(\frac{1}{p_{1}}-{\color[rgb]{0,0,0}{\frac{1}{2}}}\right)\frac{\alpha}{\alpha+\delta},\quad\delta:=\frac{1}{p_{1}}-\frac{1}{p_{2}}.$ (5.9)
@@
$\begin{split}\big\|u-\bar{{\mathcal{S}}}^{V}_{n}u\big\|_{L_{2}({\mathbb{R}}^{\infty},V;\gamma)}\ \leq\ \begin{cases}C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+1+\varepsilon}\quad&{\rm if}\ \alpha<{\color[rgb]{0,0,0}{1/p_{2}-1/2}},\\ C_{\varepsilon}n^{-\alpha}(\log n)^{\alpha+2+\varepsilon}\quad&{\rm if}\ \alpha={\color[rgb]{0,0,0}{1/p_{2}-1/2}},\\ Cn^{-\beta}(\log n)^{\beta+\alpha/(\alpha-\beta)}\ \ \quad&{\rm if}\ \alpha>{\color[rgb]{0,0,0}{1/p_{2}-1/2}},\end{cases}\end{split}$
where the positive constants $C$ depends only on $\alpha,q_{1},q_{2}$ and $K$; the positive constants $C_{\varepsilon}$ depend only on $\alpha,q_{1},q_{2},K,$ and, explicitly, on $\varepsilon$.
@@
[…]
@@
$\big|\operatorname{SamplePts}{\mathcal{I}}_{\xi_{n}}\big|\ \leq\ n,$ (5.12)
@@
[…]